### v1

# LLM-Guided Open RAN: Empowering Hierarchical RAN Intelligent Control

Abstract
Recent advancements in large language models have led to a significant interest in deploying LLM-empowered algorithms for wireless communication networks. Meanwhile, open radio access network (O-RAN) techniques offer unprecedented flexibility, with the (non-RT RIC) and (near-RT RIC) components enabling intelligent resource management across different time scales. In this paper, we propose the LLM empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs. This framework integrates LLMs with reinforcement learning (RL) for efficient network resource management. In this framework, LLMs-empowered non-RT RICs provide strategic guidance and high-level policies based on environmental context. Concurrently, RL-empowered near-RT RICs perform low-latency tasks based on strategic guidance and local near-RT observation. We evaluate the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting. Simulation results demonstrate that the proposed framework achieves superior performance. Finally, we discuss the key future challenges in applying LLMs to O-RAN.
Index Terms:
Open radio access network, large language model, reinforcement learning
I Introduction
The O-RAN has recently gained significant attention for its ability to prompt interoperability and flexibility. In O-RAN, the next generation node B (gNB) is functionally split into three components: the radio unit (RU), the distributed unit (DU), and the central unit (CU). The different components of the networks can be developed and supplied by various vendors through open standards and interfaces. This openness enables a competitive ecosystem, driving innovation and reducing costs. Moreover, by introducing a modular, software-driven architecture, O-RAN transforms the traditional hardware-centric approach to virtual RAN [1].
A core component of O-RAN is the RICs, which provides advanced control and optimization capabilities. The RICs include: 1) the non-RT RIC, which integrates with the network orchestrator and operates on a time scale larger than 1 s, and 2) the near-RT RIC, which manages control loops with RAN nodes on a time scale between $10$ ms and $1$ s. Those RICs are capable of aggregating key performance measurements (KPMs) to monitor the status of network infrastructure (e.g., number of users, load, throughput, resource utilization). Specifically, the near-RT RIC connects to the RAN components through E2 interface. The non-RT RIC interfaces with RAN components through O1 interface, and based on the available KPMs, it facilitates a non-real-time control loop and supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via A1 interface. The third-party applications, known as xApp and rApp, are deployed within near-RT RIC and non-RT RIC, respectively. These applications enable advanced functionalities, such as enhanced radio resource management, improved energy efficiency, and predictive maintenance [2].
Significant researches have been conducted to enhance RICs functionalities [3, 4, 5, 6, 7, 8]. Specifically, near-RT RIC, equipped with xApps, manages a smaller number of devices and enable low-latency optimizations. These optimizations include radio card switching to improve energy efficiency [3], flow-split distribution, congestion control, and scheduling to optimize network utility [4]. In contrast, the non-RT RIC, equipped with rApps, is responsible for managing a larger number of devices and performing high-level, long-term network management [5, 6, 7, 8]. For example, the orchestRAN framework in [5] automatically computes the optimal set of data-driven algorithms and determines their execution locations. Furthermore, cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides near-RT RIC with high-level guidance, including network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then leverages real-time information to refine and implement these policies, enhancing dynamic network performance. Despite these advancements, two critical challenges remain: 1) Traditional machine learning (ML) techniques struggle to process the diverse and complex information in O-RAN. 2) Ensuring high-quality policy generation in the non-RT RIC across diverse scenarios and improving efficient cooperation between non-RT RIC and near-RT RIC are challenging.
Meanwhile, the LLMs, such as OpenAI’s ChatGPT, Google’s Gemini, and Meta AI’s Llama, are advanced ML models, designed to understand and generate human language. Trained on vast and diverse datasets, LLMs excel in tasks such as text generation, translation, summarization, and question answering. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Telecom operators can leverage LLMs to enhance network management and improve overall service quality. Additionally, LLMs support the development of intelligent applications for network operations, including predictive maintenance, anomaly detection, and performance optimization. The LLMs have opened new routes for wireless network design and operation, offering the potential for significant performance improvements.
Recent studies have applied LLM-based technologies to wireless network for resource allocation[9, 10, 11], network deployment[12, 13] and packet analysis [14]. Specifically, [9] develope a LLM empowered-xApp to allocate physical resource block (PRB) for 5G O-RAN. However it lacks a comparison with state of art resource allocation algorithms and does not analyze the inference latency of LLMs. [10] and [11] utilize Meta AI’s Llama to derive the power allocation policy directly. Additionally, [12] deploys a combined LLM and conventional neural network (CNN) model as a RL agent for base station placement. Furthermore, the GPT model has been utilized to formulate mathematical models and generate solution codes for base station placement [13]. Although these applications show the potential of LLMs in wireless networks, several challenges remain. First, the high computational complexity hinders near-RT decision-making. Additionally, without fine-tuning for domain-specific tasks, their performance often remains comparable to traditional algorithms.
Motivated by these opportunities and challenges, we propose the LLM-hRIC framework for O-RAN. This framework deploys LLMs-empowered non-RT RIC to provide guidance for the RL-empowered near-RT RIC. The near-RT RIC utilizes this guidance along with near-RT environment information to perform near-RT decision-making. The contributions of this work are summarized as follows:
  • LLM-empowered non-RT RIC: We propose a LLM-empowered non-RT RIC that collects and processes comprehensive environmental data to provide strategic guidance to the near-RT RIC. This approach improves data integration and long-term decision-making in the O-RAN framework.
  • RL-empowered near-RT RIC: We introduce a RL-empowered near-RT RIC that utilizes the guidance from non-RT RIC and local observations to make near-RT decisions. To accelerate convergence and improve overall system performance, we deploy a joint training method that smoothly mixes non-RT RIC’s guidance and near-RT RIC’s policy.
  • Performance Evaluation: We evaluate the proposed framework in IAB networks. The results show significant performance improvements compared to baseline methods, demonstrating its ability to manage the complex information in O-RAN systems.
II LLM-empowerd hierarchical RIC Framework
In this section, we propose the LLM-empowerd hierarchical RIC (LLM-hRIC) framework for the resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages on generalization across various environments with minimal language data, it is used for the non-RT RIC to provide a guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Proposed LLM-hRIC framework.
II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, is responsible for utilizing those information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enables to form a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide a high-quality strategic guidance to the near-RT RIC.
The process of the LLM-empowered non-RT RIC can be summarized as follows.
  • 1)
    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images), and integrates the collected data by structuring it into an unified format.
  • 2)
    Environment Understanding and Actionable Insights Generation: The LLM analyze and understand the context based on the integrated multimodal data, and generate a strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control). The mathematical formulation of the guidance generation process is given as
    $\textbf{y}=\text{LLM}_{\theta}(x_{1},x_{2},\cdots,x_{N}),$ (1)
    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$, and $\theta$ denotes the parameters of LLM.
  • 3)
    Guidance transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via A1 interface to be used for the near-RT decision-making process.
II-B RL-Empowered near-RT RIC
The near-RT RIC operates as a politic layer of O-RAN, managing and optimizing network operations in near-RT. To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface. Specifically, the optimal action $\boldsymbol{a}$ can be determined as
$\boldsymbol{a}=\text{RL}_{\phi}(\textbf{y},\boldsymbol{o}),$ (2)
where $\phi$ represents the parameters of the RL model.
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the RL-based xApps need to be designed to process information efficiently and make a decision fast. From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments. This can enable the RL-based xApps to make a decision with the large-scale view on the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
III Use Case: Power Allocation in IAB Network
In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework.
III-A Power Allocation for IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as =$\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users, $\mathcal{K}$, follow a Gaussian-Markov mobility model and always connect to the nearest SBS. The SBS allocates fixed power to their connected user.
The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. The large-scale fading gain of LoS and NLoS components are given by $L^{\text{LoS}}$ and $L^{\text{NLoS}}$, respectively. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
In this use case, the total channel bandwidth $W$ for each MBS is allocated between the backhaul and access links in proportion $\alpha$, where $\alpha\in[0,1]$, i.e., the fraction of bandwidth allocated to the backhaul link is $W\alpha$ and the remaining bandwidth is assigned to the access links. At the $t$-th time slot, the achievable rate of the access link between the $k$-th user and the $n$-th SBS and that of the backhaul link between the $m$-th MBS and the $n$-th SBS are denoted by $R_{m,n,k}^{\text{a}}[t]$ and $R_{m,n}^{\text{b}}[t]$, respectively, and then the total throughput becomes $\mathcal{T}=\sum_{t=1}^{T}\sum_{m=1}^{M}\sum_{n\in\mathcal{N}_{m}}\min\left(R_{m,n}^{\text{b}}[t],\sum_{k\in\mathcal{K}_{m,n}}R_{m,n,k}^{\text{a}}[t]\right)$. The objective is to maximize the total throughput $\mathcal{T}$ by optimizing the power allocation of MBSs to each SBSs under the total power constraint on the MBS, given by $\sum_{n\in\mathcal{N}_{m}}P_{m,n}[t]\leq P_{m}^{\text{max}},\forall m,t$, where $P_{m}^{\text{max}}$ is the maximum available power of the $m$-th MBS.
III-B LLM-hRIC Framework for Power Allocation
In this section, the proposed LLM-hRIC framework is deployed for power allocation in the IAB networks. Each MBS connects to a near-RT RIC, while a centralized non-RT RIC collects information from all MBSs and provides strategic power allocation guidance to the near-RT RICs.
III-B1 LLM-Empowered non-RT RIC
In this framework, the LLM is deployed in the non-RT RIC as a rApp. A structured prompt guides the LLM in generating the power allocation guidance for near-RT RIC (e.g., initial power allocation policy). This prompt includes the following components:
  • Objective Function: Defines the goal of power allocation optimization, i.e., to maximize the total throughput.
  • Environment Description: Details the number of MBSs, SBSs, the maximum power of each MBS, etc,.
  • Channel Information: Provides the average channel gain between MBSs and SBSs.
  • Interference Information: Specifies interference sources and their average interference channel gains.
  • Constraints: Constraints for power allocation policy and the guidance format.
The structure of the prompt is as follows:
Prompt Structure You are an expert in wireless communications for resource allocation. Your objective is <Objective function >. Here is the system description: <Environment information >. Input Format: For each MBS, you are provided with the following input data for its connected SBSs. Each SBS is represented as a list: [Average channel gain, number of connected users, average expected data rate of connected users (Mb/s), and interference: [interference source (MBS,SBS), average interference channel gain] ]. Constraints: Ensure the total power allocation across SBSs for each MBS sums to 1. For each MBS, output the normalized power allocation ratios as a list of six values corresponding to its six SBSs. MBSX: [value1, value2, value3, value4, value5, value6].
Due to the time-varying nature of small-scale fading, the non-RT RIC uses average values of small-scale fading parameters to compute metrics like channel gain, expected data rate, and interference. These average values effectively represent the channel’s state in future time slots. The LLM generates the initial power allocation policy as a guidance, denoted as $\boldsymbol{p}^{o}_{m}$, for the $m$-th MBS. This guidance is updated every $T_{0}$ time slots to adapt to network environment change.
III-B2 RL-Empowered near-RT RIC
In the IAB networks, each MBS is associated RL-based xApps for near-RT resource management. The markov decision process (MDP) modeled as following components:
  • State Space: We define the system state space of $m$-th MBS as $\mathcal{S}_{m}=\{\boldsymbol{h}_{m},\boldsymbol{n}_{m},\boldsymbol{R}_{m},\boldsymbol{p}^{o}_{m}\}$, where $\boldsymbol{h}_{m}$ is the channel gain between MBS and its connected SBSs, $\boldsymbol{n}_{m}$ represents the number of associated users per connected SBS, $\boldsymbol{R}_{m}$ is the average expected data rate of connected users per SBS, and $\boldsymbol{p}^{o}_{m}$ is the initial power allocation policy provided by the non-RT RIC.
  • Action Space: Each MBS allocates power to its connected SBSs, denoted as $\boldsymbol{p}_{m}=(p_{m,1},\cdots,p_{m,N})$, where $p_{m,n}\in[0,1]$ is the power allocation ratio for each connected SBS.
  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS, and $r^{\text{g}}$ is the total throughput of all MBSs.
III-C Cooperative Training
The RL agent in the near-RT RIC need to be trained to effectively utilize the guidance provide by LLM. Traditional RL typically relies on random noise for action exploration during the initial training stages. To enhance exploration efficiency, the initial actions are instead guided by the LLM, e.g., initial power allocation policy. And then the RL action will mix with LLM guidance gradually. The training is divided into the following three phases:
  • Phase 1: The power allocation policy $\boldsymbol{p}_{m}$ is generated by adding noise to the policy $\boldsymbol{p}_{m}^{o}$ provided by the LLM, facilitating exploration around the LLM’s recommendations.
  • Phase 2: The power allocation policy $\boldsymbol{p}_{m}$ is obtained by linearly blending the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$ using a decay weighted $w$: $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where $w$ gradually decays from $1$ to 00 during this phase.
  • Phase 3: The power allocation policy $\boldsymbol{p}_{m}$ is determined by the RL agent.
In each training stage, the power allocation policy interact with environment, and the resulting MDP experiences are collected to continuously train the RL agent.
III-D Experimental Results
In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 1s). Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 100ms). To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 2: Training curve for all methods when $\alpha=0.5$.
Figure 2 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$. The results show that the total throughput increases steadily as training progresses and gradually converges as the models learn effective policies. Notably, the proposed LLM-hRIC framework achieves higher total throughput compared to the baseline methods. This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance.
Figure 3: Effect of backhaul-access bandwidth partitioning ratio $\alpha$ on total throughput.
Figure 3 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput. The result shows that the throughput initially increases and then decreases as the allocation ratio $\alpha$ increases. This is because the performance bottleneck exists at the backhaul link for a low $\alpha$, while it is primarily in the access link for large $\alpha$. The proposed LLM-hRIC framework consistently outperforms baseline methods. This is due to the high-quality initial policies, generated by LLM, which enhances exploration efficiency and facilitates better policy discovery by RL. The LLM leverage detailed, and global information to improve coordination among MBSs, resulting in a more efficient resource allocation and higher overall network performance.
IV Future Challenges
In this section, we discusses the future challenges of LLMs in O-RAN. As shown in Fig. 4, key challenges include leveraging multi-modal data, fine-tuning of LLMs, enabling systematic collaboration between RICs, and achieving timely inference. Each topic is discussed in detailed below.
Figure 4: Future challenges.
IV-A Leveraging Multi-Model Data
Multi-modal data (e.g. text, signals, images, and videos) can be used for better understanding on the complex network environments. The multi-modal LLMs are capable of reasoning across various data types, providing richer contextual insights and enabling more informed decision-making.
Meanwhile, the zero-shot or few shot learning methods (e.g. retrieval augmentd generation (RAG) and prompt engineering) have been explored to adapt LLMs to new environments. However, several open questions remain. First, current prompt engineering methods generally lack support for multi-modal inputs in wireless communication domain, making it critical to develop methods that effectively handle such data. Moreover, the performance of RAG is highly depend on retrieval engine. Designing an effective multi-modal RAG engine in wireless communication domain is key challenge for improving LLM reasoning.
IV-B Fine-Tuning of LLMs
While current LLMs have been trained well for general domain, they are not specifically optimized for wireless communication tasks, which limits their performance in this area. Although, prompt engineering can guide LLMs to some extent, limitations in context window size and the dynamic nature of wireless environment hinder real time reasoning.
To address this issue, fine-tuning LLMs using offline, domain-specific data is essential. Fine-tuning LLMs can enhance decision-making relevance and reliability. Nonetheless, two major open challenges remain: 1) how to collect or effectively utilize limited wireless communication data, and 2) how to effectively generate high quality instruction data for fine-tuning?
IV-C Systematic Collaboration between RICs
Due to the lack of domain specific fine-tuning, LLMs-generated policies are not always optimal. Enhancing them with lightweight AI techniques is critical, especially for near-RT applications where inference speed is paramount. To fully exploit the strength of both LLMs (for long time, global optimization) and lightweight AI (for real-time responsiveness), a cooperative framework must be established. This framework should enable smooth interaction between the LLMs and lightweight AI models, supporting adaptive and robust decision-making under dynamic conditions.
Several challenges need to be addressed. First, LLMs operate on large time scale, while network environment change rapidly in near-RT. Bridging this time gap to make training and prediction work well is not easy. Second, due to the limited capacity of lightweight models, selecting the most relevant features for specific tasks is critical for performance enhancement.
IV-D Timely Inference
Modern LLMs, such as GPT-4 with 1.7 trillion parameters, demand significant computational resources and are often deployed on cloud-based servers. This dependency introduces substantial latency, making deploying LLM model in near-RT RIC as a xApp challenging.
To address this challenge, technologies such as model pruning, quantization and knowledge distillation are crucial for reduce the model size. Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead without significantly affecting performance. In addition, the model pruning and knowledge distillation train a smaller, optimized versions of LLMs. The different methods can compress the model with different ratio. However Compression methods typically trade off some model performance for reduced latency. Balancing model size, inference speed, and accuracy is a critical challenge.
V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM is deployed as rApp in the non-RT RIC to provide strategic guidance, while RL is implemented as xApp in the near-RT RIC to facilitate near-RT decision-making. The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems.

### v2

@@
-LLM-Guided Open RAN: Empowering Hierarchical RAN Intelligent Control
+LLM-hRIC: LLM-empowered Hierarchical RAN Intelligent Control for O-RAN
 Abstract
-Recent advancements in large language models have led to a significant interest in deploying LLM-empowered algorithms for wireless communication networks.
-Meanwhile, open radio access network (O-RAN) techniques offer unprecedented flexibility, with the (non-RT RIC) and (near-RT RIC) components enabling intelligent resource management across different time scales.
-In this paper, we propose the LLM empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs.
-This framework integrates LLMs with reinforcement learning (RL) for efficient network resource management.
-In this framework, LLMs-empowered non-RT RICs provide strategic guidance and high-level policies based on environmental context.
-Concurrently, RL-empowered near-RT RICs perform low-latency tasks based on strategic guidance and local near-RT observation.
-We evaluate the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting.
-Simulation results demonstrate that the proposed framework achieves superior performance.
-Finally, we discuss the key future challenges in applying LLMs to O-RAN.
+Despite recent advances in applying large language models and machine learning (ML) techniques to open radio access network (O-RAN), critical challenges remain, such as insufficient cooperation between RICs, high computational demands hindering real-time decisions, and the lack of domain-specific fine-tuning.
+Therefore, this article introduces the LLM-empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs in O-RAN.
+The LLM-empowered non-real-time (non-RT RIC) acts as a guider, offering a strategic guidance to the near-real-time (near-RT RIC) using global network information.
+The RL-empowered near-RT RIC acts as an implementer, combining this guidance with local real-time data to make near-RT decisions.
+We evaluate the feasibility and performance of the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting, and finally, discuss the open challenges of the LLM-hRIC framework for O-RAN.
 Index Terms:
-Open radio access network, large language model, reinforcement learning
+Open radio access network, large language model, RAN intelligent control, reinforcement learning
 I Introduction
@@
 The RICs include: 1) the non-RT RIC, which integrates with the network orchestrator and operates on a time scale larger than 1 s, and 2) the near-RT RIC, which manages control loops with RAN nodes on a time scale between $10$ ms and $1$ s.
-Those RICs are capable of aggregating key performance measurements (KPMs) to monitor the status of network infrastructure (e.g., number of users, load, throughput, resource utilization).
-Specifically, the near-RT RIC connects to the RAN components through E2 interface.
-The non-RT RIC interfaces with RAN components through O1 interface, and based on the available KPMs, it facilitates a non-real-time control loop and supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via A1 interface.
-The third-party applications, known as xApp and rApp, are deployed within near-RT RIC and non-RT RIC, respectively.
-These applications enable advanced functionalities, such as enhanced radio resource management, improved energy efficiency, and predictive maintenance [2].
-Significant researches have been conducted to enhance RICs functionalities [3, 4, 5, 6, 7, 8].
-Specifically, near-RT RIC, equipped with xApps, manages a smaller number of devices and enable low-latency optimizations.
-These optimizations include radio card switching to improve energy efficiency [3], flow-split distribution, congestion control, and scheduling to optimize network utility [4].
-In contrast, the non-RT RIC, equipped with rApps, is responsible for managing a larger number of devices and performing high-level, long-term network management [5, 6, 7, 8].
-For example, the orchestRAN framework in [5] automatically computes the optimal set of data-driven algorithms and determines their execution locations.
-Furthermore, cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8].
-Specifically, the non-RT RIC provides near-RT RIC with high-level guidance, including network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8].
-The near-RT RIC then leverages real-time information to refine and implement these policies, enhancing dynamic network performance.
-Despite these advancements, two critical challenges remain: 1) Traditional machine learning (ML) techniques struggle to process the diverse and complex information in O-RAN.
-2) Ensuring high-quality policy generation in the non-RT RIC across diverse scenarios and improving efficient cooperation between non-RT RIC and near-RT RIC are challenging.
-Meanwhile, the LLMs, such as OpenAI’s ChatGPT, Google’s Gemini, and Meta AI’s Llama, are advanced ML models, designed to understand and generate human language.
-Trained on vast and diverse datasets, LLMs excel in tasks such as text generation, translation, summarization, and question answering.
+The non-RT RIC and the near-RT RIC are capable of obtaining key performance measurements (KPMs) via O1 and E2 interfaces, respectively, to monitor the status of network infrastructure (e.g., number of users, load, throughput, and resource utilization).
+The non-RT RIC also supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via the A1 interface.
+Specific RIC functionalities, such as mobility management and network slicing, are deployed within near-RT RIC and non-RT RIC in the form of the third-party applications, known as xApp and rApp, respectively [2].
+Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8].
+Specifically, the near-RT RICs has been developed to manage a smaller number of devices and optimize on a small-time scale, such as the radio card switching for energy efficiency [3] and the flow-split distribution, congestion control, and scheduling for higher network utility [4].
+The non-RT RIC has been developed to manage a larger number of devices and perform high-level and long-term network management, such as the orchestRAN[5], which automatically computes the optimal set of algorithms and their execution location.
+Furthermore, the cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8].
+Specifically, the non-RT RIC provides the high-level guidance to near-RT RIC, including the network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8].
+The near-RT RIC then refines and implements these policies based on real-time information to enhance the network performance.
+Meanwhile, the LLMs have opened new routes for wireless network design and operation including O-RAN.
+The LLMs are advanced ML models, designed to understand and generate human language.
 Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains.
 In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making.
-Telecom operators can leverage LLMs to enhance network management and improve overall service quality.
-Additionally, LLMs support the development of intelligent applications for network operations, including predictive maintenance, anomaly detection, and performance optimization.
-The LLMs have opened new routes for wireless network design and operation, offering the potential for significant performance improvements.
-Recent studies have applied LLM-based technologies to wireless network for resource allocation[9, 10, 11], network deployment[12, 13] and packet analysis [14].
-Specifically, [9] develope a LLM empowered-xApp to allocate physical resource block (PRB) for 5G O-RAN.
-However it lacks a comparison with state of art resource allocation algorithms and does not analyze the inference latency of LLMs.
-[10] and [11] utilize Meta AI’s Llama to derive the power allocation policy directly.
-Additionally, [12] deploys a combined LLM and conventional neural network (CNN) model as a RL agent for base station placement.
-Furthermore, the GPT model has been utilized to formulate mathematical models and generate solution codes for base station placement [13].
-Although these applications show the potential of LLMs in wireless networks, several challenges remain.
-First, the high computational complexity hinders near-RT decision-making.
-Additionally, without fine-tuning for domain-specific tasks, their performance often remains comparable to traditional algorithms.
-Motivated by these opportunities and challenges, we propose the LLM-hRIC framework for O-RAN.
-This framework deploys LLMs-empowered non-RT RIC to provide guidance for the RL-empowered near-RT RIC.
-The near-RT RIC utilizes this guidance along with near-RT environment information to perform near-RT decision-making.
-The contributions of this work are summarized as follows:
-  • LLM-empowered non-RT RIC: We propose a LLM-empowered non-RT RIC that collects and processes comprehensive environmental data to provide strategic guidance to the near-RT RIC.
-This approach improves data integration and long-term decision-making in the O-RAN framework.
-  • RL-empowered near-RT RIC: We introduce a RL-empowered near-RT RIC that utilizes the guidance from non-RT RIC and local observations to make near-RT decisions.
-To accelerate convergence and improve overall system performance, we deploy a joint training method that smoothly mixes non-RT RIC’s guidance and near-RT RIC’s policy.
-  • Performance Evaluation: We evaluate the proposed framework in IAB networks.
-The results show significant performance improvements compared to baseline methods, demonstrating its ability to manage the complex information in O-RAN systems.
+Recent studies have applied LLM-based technologies to wireless network for the resource allocation[9, 10], base station placement[11, 12], and packet analysis [13].
+Recently, the LLM-based xApp to allocate physical resource block (PRB) for O-RAN has also been introduced[14].
+Despite these advancements in RICs and LLM-based techniques for wireless networks, critical challenges still remain as follows: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
+Motivated by these opportunities and challenges, we propose the novel framework for O-RAN, which is the LLM-hRIC.
+In this framework, there are two RICs: 1) the LLMs-empowered non-RT RIC and 2) the reinforcement learning (RL)-empowered near-RT RIC.
+The LLMs-empowered non-RT RIC works as a guider to the near-RT RIC.
+Based on the objective of the near-RT RIC, it provides strategic guidance for the RL-empowered near-RT RIC based on the global view of the networks (e.g., long-term channel information, network nodes state, users mobility model, and users distribution).
+The RL-empowered near-RT RIC works as an implementer, which utilizes this guidance along with the local view of the networks (e.g., real-time channel quality and local network node state) to perform near-RT decision-making.
+The key contribution of the paper can be summarized as follows.
+  • We propose the novel LLM-hRIC framework for O-RAN, which can show how the LLM can be utilized to enhance the performance of RICs in O-RAN.
+This article focuses on addressing the limitation of LLM in making near-RT decisions and paves a way to utilize it in a reliable way in O-RAN environments.
+  • We introduce the three sequential phases for training of the RL-empowered near-RT RIC, of each phase utilizes the guidance of the non-RT RIC in a different way to accelerate convergence as well as improve the system performance.
+  • We quantify the performance of the LLM-hRIC for the power allocation in the IAB networks as a use case. This can demonstrate the ability of LLM-hRIC to manage the complex information in O-RAN.
 II LLM-empowerd hierarchical RIC Framework
-In this section, we propose the LLM-empowerd hierarchical RIC (LLM-hRIC) framework for the resource management in O-RAN as shown in Fig.
+In this section, we introduce the LLM-empowered hierarchical RIC (LLM-hRIC) framework for resource management in O-RAN as shown in Fig.
 1.
 The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN.
-As the LLM has the advantages on generalization across various environments with minimal language data, it is used for the non-RT RIC to provide a guidance on the strategies to the near-RT RIC from the large-scale view of the networks.
+As the LLM has the advantages of generalization across various environments with minimal language data, it is used for the non-RT RIC to provide guidance on the strategies to the near-RT RIC from the large-scale view of the networks.
 The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT.
 More details on the LLM-hRIC framework are provided in the following subsections.
-Figure 1: Proposed LLM-hRIC framework.
+Figure 1: Structure of the proposed LLM-hRIC framework.
 II-A LLM-Empowered non-RT RIC
 The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface.
-The rApps, deployed for the non-RT RIC, is responsible for utilizing those information to generate strategic guidance for the near-RT RIC.
-By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enables to form a comprehensive understanding of the current network environments.
+The rApps, deployed for the non-RT RIC, are responsible for utilizing that information to generate strategic guidance for the near-RT RIC.
+By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enabling to form of a comprehensive understanding of the current network environments.
 Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference.
-These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide a high-quality strategic guidance to the near-RT RIC.
+These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide high-quality strategic guidance to the near-RT RIC.
 The process of the LLM-empowered non-RT RIC can be summarized as follows.
   • 1)
-    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images), and integrates the collected data by structuring it into an unified format.
+    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images) and integrates the collected data by structuring it into a unified format.
   • 2)
-    Environment Understanding and Actionable Insights Generation: The LLM analyze and understand the context based on the integrated multimodal data, and generate a strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control).
+    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control).
 The mathematical formulation of the guidance generation process is given as
     $\textbf{y}=\text{LLM}_{\theta}(x_{1},x_{2},\cdots,x_{N}),$ (1)
-    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$, and $\theta$ denotes the parameters of LLM.
+    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$ and $\theta$ denotes the parameters of LLM.
   • 3)
-    Guidance transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via A1 interface to be used for the near-RT decision-making process.
+    Guidance Transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via the A1 interface, used for the near-RT decision-making process.
 II-B RL-Empowered near-RT RIC
-The near-RT RIC operates as a politic layer of O-RAN, managing and optimizing network operations in near-RT.
+The near-RT RIC operates as a political layer of O-RAN, managing and optimizing network operations in near-RT.
 To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface.
@@
 From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments.
-This can enable the RL-based xApps to make a decision with the large-scale view on the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
+This can enable the RL-based xApps to decide with a large-scale view of the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
 III Use Case: Power Allocation in IAB Network
 In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework.
-III-A Power Allocation for IAB Networks
+III-A Power Allocation in IAB Networks
 The IAB networks as illustrated in Fig.
@@
 Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS.
-A set of mobile users, $\mathcal{K}$, follow a Gaussian-Markov mobility model and always connect to the nearest SBS.
+A set of mobile users $\mathcal{K}_{m,n}$, connected with the $n$-th SBS of the $m$-th MBS, follow a Gaussian-Markov mobility model.
 The SBS allocates fixed power to their connected user.
@@
 Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant.
-The large-scale fading gain of LoS and NLoS components are given by $L^{\text{LoS}}$ and $L^{\text{NLoS}}$, respectively.
 Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
@@
   • Interference Information: Specifies interference sources and their average interference channel gains.
-  • Constraints: Constraints for power allocation policy and the guidance format.
-The structure of the prompt is as follows:
+  • Constraints: Details the constraints on power allocation policy and the guidance format.
+The structure of the prompt is as Fig. 2.
 Prompt Structure You are an expert in wireless communications for resource allocation.
@@
 MBSX: [value1, value2, value3, value4, value5, value6].
+Figure 2: The Prompt Structure.
 Due to the time-varying nature of small-scale fading, the non-RT RIC uses average values of small-scale fading parameters to compute metrics like channel gain, expected data rate, and interference.
 These average values effectively represent the channel’s state in future time slots.
-The LLM generates the initial power allocation policy as a guidance, denoted as $\boldsymbol{p}^{o}_{m}$, for the $m$-th MBS.
-This guidance is updated every $T_{0}$ time slots to adapt to network environment change.
 III-B2 RL-Empowered near-RT RIC
@@
   • Action Space: Each MBS allocates power to its connected SBSs, denoted as $\boldsymbol{p}_{m}=(p_{m,1},\cdots,p_{m,N})$, where $p_{m,n}\in[0,1]$ is the power allocation ratio for each connected SBS.
-  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS, and $r^{\text{g}}$ is the total throughput of all MBSs.
-III-C Cooperative Training
-The RL agent in the near-RT RIC need to be trained to effectively utilize the guidance provide by LLM.
-Traditional RL typically relies on random noise for action exploration during the initial training stages.
-To enhance exploration efficiency, the initial actions are instead guided by the LLM, e.g., initial power allocation policy.
-And then the RL action will mix with LLM guidance gradually.
-The training is divided into the following three phases:
-  • Phase 1: The power allocation policy $\boldsymbol{p}_{m}$ is generated by adding noise to the policy $\boldsymbol{p}_{m}^{o}$ provided by the LLM, facilitating exploration around the LLM’s recommendations.
-  • Phase 2: The power allocation policy $\boldsymbol{p}_{m}$ is obtained by linearly blending the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$ using a decay weighted $w$: $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where $w$ gradually decays from $1$ to 00 during this phase.
-  • Phase 3: The power allocation policy $\boldsymbol{p}_{m}$ is determined by the RL agent.
-In each training stage, the power allocation policy interact with environment, and the resulting MDP experiences are collected to continuously train the RL agent.
+  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS and $r^{\text{g}}$ is the total throughput of all MBSs.
+III-C LLM-assisted Three Phases Training for RL
+The RL agent in the near-RT RIC must be trained to effectively utilize the guidance provided by LLM.
+Traditional RL typically relies on a random policy for action exploration during the initial stage of training.
+To improve the exploration efficiency, this random policy is replaced with the LLM guidance, such as an initial power allocation policy and injected action noise.
+The policy is gradually formed by blending the actions, generated by the RL agent with the LLM guidance, until the agent is capable of making decisions independently.
+Specifically, the training can be processed by the following three sequential phases:
+  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$, to encourage exploration in the vicinity of the LLM’s guidance.
+  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 00 over time.
+  • Phase 3 (Self-directed Decision): The power allocation policy $\boldsymbol{p}_{m}$ is determined solely by the RL agent, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{d}$.
+In each training stage, the power allocation policy interacts with environment, and the resulting MDP experiences are collected to train the RL agent.
 III-D Experimental Results
-In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 1s).
-Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 100ms).
+In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 2s).
+Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 200ms).
 To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
-Figure 2: Training curve for all methods when $\alpha=0.5$.
+Figure 3: Training curve for proposed and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
 Figure 3 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$.
@@
 This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance.
-Figure 3: Effect of backhaul-access bandwidth partitioning ratio $\alpha$ on total throughput.
+Figure 4: Average total throughput of the proposed LLM-hRIC and baseline methods according to backhaul-access bandwidth partitioning ratio $\alpha$ over 50 test examples.
 Figure 4 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput.
@@
 IV Future Challenges
-In this section, we discusses the future challenges of LLMs in O-RAN.
+In this section, we discusses the future challenges of LLM-hRIC for O-RAN.
 As shown in Fig.
-4, key challenges include leveraging multi-modal data, fine-tuning of LLMs, enabling systematic collaboration between RICs, and achieving timely inference.
+5, key challenges specific to LLM-hRIC include leveraging multi-modal data, fine-tuning LLMs for guidance generation, enabling systematic collaboration between RICs in LLM-hRIC, and achieving timely inference.
 Each topic is discussed in detailed below.
-Figure 4: Future challenges.
-IV-A Leveraging Multi-Model Data
-Multi-modal data (e.g.
-text, signals, images, and videos) can be used for better understanding on the complex network environments.
-The multi-modal LLMs are capable of reasoning across various data types, providing richer contextual insights and enabling more informed decision-making.
-Meanwhile, the zero-shot or few shot learning methods (e.g.
-retrieval augmentd generation (RAG) and prompt engineering) have been explored to adapt LLMs to new environments.
-However, several open questions remain.
-First, current prompt engineering methods generally lack support for multi-modal inputs in wireless communication domain, making it critical to develop methods that effectively handle such data.
-Moreover, the performance of RAG is highly depend on retrieval engine.
-Designing an effective multi-modal RAG engine in wireless communication domain is key challenge for improving LLM reasoning.
+Figure 5: Future challenges of the LLM-hRIC for O-RAN.
+IV-A Leveraging Multi-Model Data in LLM-hRIC
+In the LLM-hRIC, the multi-modal LLMs can reason across various data types.
+Multi-modal data (e.g., text, signals, images, and videos) provides detailed environment insights, enhancing the ability of the non-RT RIC to generate more accurate and context-aware guidance for near-RT RIC.
+Moreover, zero-shot and few-shot learning techniques, such as prompt engineering and retrieval augmentd generation (RAG), have been explored to enhance the performance of LLMs in new environments.
+Despite their potential, significant challenges remain: 1) prompt engineering lacks robust support for multi-modal inputs in the wireless communication domain, necessitating new methods to handle such data for generating high-quality guidance.
+2) RAG performance relies heavily on the retrieval engine.
+Designing an effective multi-modal RAG engine for LLM-hRIC in the wireless communication domain is critical to improve LLM reasoning and guidance generation.
 IV-B Fine-Tuning of LLMs
-While current LLMs have been trained well for general domain, they are not specifically optimized for wireless communication tasks, which limits their performance in this area.
-Although, prompt engineering can guide LLMs to some extent, limitations in context window size and the dynamic nature of wireless environment hinder real time reasoning.
-To address this issue, fine-tuning LLMs using offline, domain-specific data is essential.
-Fine-tuning LLMs can enhance decision-making relevance and reliability.
-Nonetheless, two major open challenges remain: 1) how to collect or effectively utilize limited wireless communication data, and 2) how to effectively generate high quality instruction data for fine-tuning?
+Current LLMs are trained for general domains but not optimized for wireless communication tasks, limiting guidance quality in LLM-hRIC.
+While prompt engineering can partially bridge this gap, its effectiveness is limited by the quality of examples and input token window size.
+Fine-tuning LLMs with offline, domain-specific data is essential to enhance guidance reliability in the non-RT RIC.
+However, two challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging.
+2) Generating high-quality instruction data for fine-tuning is essential.
+Automated or semi-automated methods for generating diverse, domain-relevant instruction data must be developed to improve the guidance quality of fine-tuned LLMs.
 IV-C Systematic Collaboration between RICs
-Due to the lack of domain specific fine-tuning, LLMs-generated policies are not always optimal.
-Enhancing them with lightweight AI techniques is critical, especially for near-RT applications where inference speed is paramount.
-To fully exploit the strength of both LLMs (for long time, global optimization) and lightweight AI (for real-time responsiveness), a cooperative framework must be established.
-This framework should enable smooth interaction between the LLMs and lightweight AI models, supporting adaptive and robust decision-making under dynamic conditions.
+In LLM-hRIC, the LLMs-empowered non-RT RIC generates guidance for multiple near-RT RICs, each responsible for diverse tasks.
+A systematic collaboration framework is needed to ensure smooth interaction between the LLMs and lightweight AI models.
 Several challenges need to be addressed.
-First, LLMs operate on large time scale, while network environment change rapidly in near-RT.
-Bridging this time gap to make training and prediction work well is not easy.
-Second, due to the limited capacity of lightweight models, selecting the most relevant features for specific tasks is critical for performance enhancement.
+1) non-RT RICs operates on a large time scale, while near-RT RICs can operate on different small time scales.
+This temporal mismatch complicates the synchronization between the guidance generation in the non-RT RIC and the real-time decision-making in the near-RT RIC.
+2) near-RT RICs perform a variety of tasks (e.g., power allocation, PRB allocation, and user association).
+Each task requires specific guidance tailored to its unique requirements, making it crucial to develop a mechanism for providing task-aware, customized guidance for different near-RT RICs.
 IV-D Timely Inference
-Modern LLMs, such as GPT-4 with 1.7 trillion parameters, demand significant computational resources and are often deployed on cloud-based servers.
-This dependency introduces substantial latency, making deploying LLM model in near-RT RIC as a xApp challenging.
-To address this challenge, technologies such as model pruning, quantization and knowledge distillation are crucial for reduce the model size.
-Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead without significantly affecting performance.
-In addition, the model pruning and knowledge distillation train a smaller, optimized versions of LLMs.
-The different methods can compress the model with different ratio.
-However Compression methods typically trade off some model performance for reduced latency.
-Balancing model size, inference speed, and accuracy is a critical challenge.
+Modern LLMs, such as GPT-4 with 1.7 trillion parameters, require significant computational resources and are often cloud-deployed, causing latency issues.
+This makes deploying an LLM in the non-RT RIC challenging, as some non-RT RIC require latency near 1s.
+Technologies like model pruning, quantization, and knowledge distillation can reduce model size.
+Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead, while pruning and distillation create smaller, optimized LLMs.
+These methods compress models at varying ratios but often trade off performance for speed.
+Balancing model size, inference speed, and accuracy is critical to ensuring efficient deployment of LLMs in the LLM-hRIC framework.
 V Conclusions
 This paper proposes the LLM-hRIC framework for O-RAN.
-In this framework, the LLM is deployed as rApp in the non-RT RIC to provide strategic guidance, while RL is implemented as xApp in the near-RT RIC to facilitate near-RT decision-making.
+In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores guidance with local observation to facilitate near-RT decision-making.
 The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management.
 Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems.
+Additionally, we identify key challenges for LLM-hRIC, as discussed in Section IV, to guide future research and development.

### v3

@@
-LLM-hRIC: LLM-empowered Hierarchical RAN Intelligent Control for O-RAN
+arXiv:2504.18062v3[cs.NI] 27 Feb 2026
 Abstract
@@
 We evaluate the feasibility and performance of the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting, and finally, discuss the open challenges of the LLM-hRIC framework for O-RAN.
-Index Terms:
-Open radio access network, large language model, RAN intelligent control, reinforcement learning
 I Introduction
@@
 Specific RIC functionalities, such as mobility management and network slicing, are deployed within near-RT RIC and non-RT RIC in the form of the third-party applications, known as xApp and rApp, respectively [2].
-Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8].
+Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8, 9].
 Specifically, the near-RT RICs has been developed to manage a smaller number of devices and optimize on a small-time scale, such as the radio card switching for energy efficiency [3] and the flow-split distribution, congestion control, and scheduling for higher network utility [4].
+In particular, deep reinforcement learning (DRL) has emerged as a key technology for intelligent xApps, enabling dynamic control of radio resource management in O-RAN [9].
 The non-RT RIC has been developed to manage a larger number of devices and perform high-level and long-term network management, such as the orchestRAN[5], which automatically computes the optimal set of algorithms and their execution location.
@@
 In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making.
-Recent studies have applied LLM-based technologies to wireless network for the resource allocation[9, 10], base station placement[11, 12], and packet analysis [13].
+Recent studies have applied LLM-based technologies to wireless network for the resource allocation[10, 11], base station placement[12], and packet analysis [13].
 Recently, the LLM-based xApp to allocate physical resource block (PRB) for O-RAN has also been introduced[14].
-Despite these advancements in RICs and LLM-based techniques for wireless networks, critical challenges still remain as follows: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
+Despite advancements in RICs and LLM-based techniques for wireless networks, a critical real-world problem remains: bridging the ‘coordination gap’ between the strategic, long-term control of the non-RT RIC and the tactical, real-time execution of the near-RT RIC.
+This gap is exacerbated by several underlying challenges: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
 Motivated by these opportunities and challenges, we propose the novel framework for O-RAN, which is the LLM-hRIC.
@@
 As the LLM has the advantages of generalization across various environments with minimal language data, it is used for the non-RT RIC to provide guidance on the strategies to the near-RT RIC from the large-scale view of the networks.
-The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT.
+The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view to make decisions in near-RT.
 More details on the LLM-hRIC framework are provided in the following subsections.
-Figure 1: Structure of the proposed LLM-hRIC framework.
+Figure 1: Structure of the proposed LLM-hRIC framework and time diagrame.
 II-A LLM-Empowered non-RT RIC
@@
 The rApps, deployed for the non-RT RIC, are responsible for utilizing that information to generate strategic guidance for the near-RT RIC.
-By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enabling to form of a comprehensive understanding of the current network environments.
+By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multi-modal data to form a comprehensive understanding of the current network environments.
 Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference.
 These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide high-quality strategic guidance to the near-RT RIC.
-The process of the LLM-empowered non-RT RIC can be summarized as follows.
+The process of the LLM-empowered non-RT RIC can be summarized as following steps.
   • 1)
-    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images) and integrates the collected data by structuring it into a unified format.
+    Data Integration: The non-RT RIC ingests multi-modal data from gNB and near-RT RIC, primarily via the O1 interface.
+These data include long-term time-series metrics (e.g., total throughput trends), unstructured textual data (e.g., fault alarm logs), topology data representing network node connectivity, and even advanced visual data such as radio maps generated from digital twins.
   • 2)
-    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control).
-The mathematical formulation of the guidance generation process is given as
-    $\textbf{y}=\text{LLM}_{\theta}(x_{1},x_{2},\cdots,x_{N}),$ (1)
-    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$ and $\theta$ denotes the parameters of LLM.
+    Input Validation and Prompt Construction: Before being formatted into a prompt, this data undergoes a validation stage to handle potentially noisy or corrupted measurements.
+The data integration process involves structuring these varied inputs into a unified, context-rich format that can be effectively processed by the LLM.
   • 3)
+    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate the text response, that includes the strategic guidance y (e.g., transmit power allocation strategies).
+  • 4)
+    Guidance Extraction and Verification: The LLM’s raw text output is validated for its conformance to predefined formats and constraints, and extracts the strategic guidance y.
+To ensure robustness, if this process fails, a fallback mechanism provides a default policy as guidance.
+  • 5)
     Guidance Transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via the A1 interface, used for the near-RT decision-making process.
 II-B RL-Empowered near-RT RIC
-The near-RT RIC operates as a political layer of O-RAN, managing and optimizing network operations in near-RT.
-To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface.
-Specifically, the optimal action $\boldsymbol{a}$ can be determined as
-$\boldsymbol{a}=\text{RL}_{\phi}(\textbf{y},\boldsymbol{o}),$ (2)
-where $\phi$ represents the parameters of the RL model.
-Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the RL-based xApps need to be designed to process information efficiently and make a decision fast.
-From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments.
-This can enable the RL-based xApps to decide with a large-scale view of the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
+Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the xApps need to be designed to process information efficiently and make fast decisions in a highly dynamic network environment.
+Hence, we employ RL because its ability to learn an adaptive policy is well-suited for reacting to the dynamic environments with real-time changes.
+Furthermore, RL is naturally designed to incorporate a strategic guidance vector as a part of its state input, fusing it with the observed local information to make refined and context-aware decisions.
+The process of the RL-empowered near-RT RIC can be summarized in the following steps:
+  • 6)
+    Data Integration: The near-RT RIC ingests the strategic guidance from the non-RT RIC (via the A1 interface) and the observed local information from the gNB (via the E2 interface).
+  • 7)
+    Policy generation: The RL-based xApp then analyzes these inputs to generate an optimal, fine-grained policy that is sent to the gNB for execution.
+This integration of global insights with real-time local information enhances the RL agent’s ability to make refined, fine-grained decisions that are both strategically sound and tactically optimal for the immediate radio conditions.
 III Use Case: Power Allocation in IAB Network
 In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework.
+This scenario was chosen as a foundational proof-of-concept since the resource allocation in multi-cell IAB networks is a well-known difficult problem to handle as it requires the network-wise global information and control.
+This problem can demonstrate our framework’s strength: using the LLM in the non-RT RIC to provide a globally-aware and coordinated guidance to local RL agents.
 III-A Power Allocation in IAB Networks
 The IAB networks as illustrated in Fig.
-1, consists of a set of macro base stations, denoted as =$\mathcal{M}=\{1,\cdots,M\}$, distributed in a network.
+1, consists of a set of macro base stations, denoted as $\mathcal{M}=\{1,\cdots,M\}$, distributed in a network.
 Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS.
@@
 These average values effectively represent the channel’s state in future time slots.
+To ensure the reliability of these inputs, a data checking step performs range validation.
+This process verifies that all metrics are physically plausible, such as ensuring that the ‘number of connected users’ is non-negative and that the ‘average channel gain’ is within a realistic range.
 III-B2 RL-Empowered near-RT RIC
@@
 Specifically, the training can be processed by the following three sequential phases:
-  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$, to encourage exploration in the vicinity of the LLM’s guidance.
-  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 00 over time.
+  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$.
+This encourages exploration in the vicinity of the LLM’s guidance, allowing the RL agent to collect high-quality experiences and learn a near-optimal initial policy.
+  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 0 over time.
+This creates a smooth transition from guided exploration to self-directed optimization.
   • Phase 3 (Self-directed Decision): The power allocation policy $\boldsymbol{p}_{m}$ is determined solely by the RL agent, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{d}$.
-In each training stage, the power allocation policy interacts with environment, and the resulting MDP experiences are collected to train the RL agent.
+In each training stage, the power allocation policy interacts with the environment, and the resulting MDP experiences are collected to train the RL agent.
+This phased approach prevents the instability seen with abrupt policy changes, allowing the agent to efficiently fine-tune its near-optimal policy, which leads to stable convergence and the highest overall throughput.
 III-D Experimental Results
-In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 2s).
-Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 200ms).
+Our simulation deploys the Llama-3.1-8B-instruct model [15] in the rApp to generate a power allocation guidance policy on a 2s timescale.
+For efficient deployment, the model is loaded with bfloat16 precision, with inference parameters of temperature =0.6=0.6 and top_p =0.9=0.9.
+Meanwhile, a deep deterministic policy gradient (DDPG) agent in the xApp provides a fine-grained power allocation policy for a MBS on a 200ms timescale.
+The DDPG agent was trained using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 256.
+[^1] Note that the demonstration on a prototype testbed can be done once O1 and A1 interfaces become sufficiently mature in open-source O-RAN platforms.
 To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
-Figure 3: Training curve for proposed and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
-Figure 3 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$.
+Figure 3: LLM and DDPG inference latency with 500 examples.
+Fig.
+3 shows the LLM and DDPG inference latency with $500$ examples.
+The LLM inference latency is less than 1.5 seconds, confirming its decision time is well within the operational budget required for the LLM-hRIC ($\leq$ 2s).
+Moreover, the DDPG inference time is less than $0.07$ ms.
+This extremely low latency is significantly faster than the 10 ms requirement for the most demanding near-real-time control.
+Figure 4: Training curve for proposed with different blending coefficient $w$ and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
+Fig.
+4 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$.
 The results show that the total throughput increases steadily as training progresses and gradually converges as the models learn effective policies.
@@
 This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance.
+Furthermore, the decaying $w$ is critical for performance.
+A fixed $w=0$ leads to training failure from an abrupt policy shift, while a fixed $w=0.9$ causes performance to stagnate by suppressing the agent’s exploration.
+In contrast, the proposed decaying $w$ ensures a smooth transition from guided exploration to self-directed optimization, which is necessary for stable convergence to the highest-performing policy.
 Figure 5: Average total throughput of the proposed LLM-hRIC and baseline methods according to backhaul-access bandwidth partitioning ratio $\alpha$ over 50 test examples.
-Figure 4 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput.
+Fig.
+5 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput.
 The result shows that the throughput initially increases and then decreases as the allocation ratio $\alpha$ increases.
@@
 As shown in Fig.
-5, key challenges specific to LLM-hRIC include leveraging multi-modal data, fine-tuning LLMs for guidance generation, enabling systematic collaboration between RICs in LLM-hRIC, and achieving timely inference.
+6, key challenges specific to LLM-hRIC include leveraging multi-modal data, fine-tuning LLMs for guidance generation, enabling systematic collaboration between RICs in LLM-hRIC, and achieving timely inference.
 Each topic is discussed in detailed below.
@@
 While prompt engineering can partially bridge this gap, its effectiveness is limited by the quality of examples and input token window size.
+Moreover, the guidance from a general-purpose LLM can sometimes be unstable, making fine-tuning important for adapting to environmental changes and ensuring reliability.
 Fine-tuning LLMs with offline, domain-specific data is essential to enhance guidance reliability in the non-RT RIC.
-However, two challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging.
+However, three challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging.
 2) Generating high-quality instruction data for fine-tuning is essential.
 Automated or semi-automated methods for generating diverse, domain-relevant instruction data must be developed to improve the guidance quality of fine-tuned LLMs.
+3) A privacy-preserving architecture should be considered such as the combination of federated learning (FL) with differential privacy (DP).
+The DP adds controlled noise to the shared model updates in FL for a mathematical guarantee that private information on the training data is not leaked.
 IV-C Systematic Collaboration between RICs
@@
 This paper proposes the LLM-hRIC framework for O-RAN.
-In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores guidance with local observation to facilitate near-RT decision-making.
-The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management.
+In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores the guidance with local observation to facilitate near-RT decision-making.
+The superior performance demonstrates the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management.
 Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems.
-Additionally, we identify key challenges for LLM-hRIC, as discussed in Section IV, to guide future research and development.
+For instance, the framework is extensible to advanced paradigms like the intent-based networking, where the LLM can translate an administrator’s high-level and multi-objective intent into strategic guidance.
+This guidance is then executed by RL agents with real-time and tactical adaptations, showcasing a powerful synergy for future network automation.
+Additionally, we identify key challenges for LLM-hRIC, as discussed in Section ‘Future Challenges’, to guide future research and development.